# Unforgettable Lessons from Forgettable Images: Intra-Class Memorability Matters in Computer Vision

### v4

@@ A.4 Performance of Models Using Memorability Predictors from the Previous Works
Figure A4: Performance comparison in Top-1 Accuracy between image classification models trained on HM, MM, and LM images using different memorability predictors. (a) shows results using our ICMscore predictor from Sec. 4.1, while (b) presents results from the regressor proposed in [5]. Within each subplot, from the top to the bottom, top-1 accuracy is reported for models trained on LM images (blue), HM images (green), and MM images (yellow). The chance is 20% for the 5-way classification (dash lines produced by the random model in red). The error bar is the standard error calculated over 5 independent runs with different random seeds.
((c))
((d))
@@
Figure A5: Continual learning performance of AI models using HM, MM, and LM images for replays (a) shows results using our ICMscore predictor from Sec. 4.1, while (b) presents results from the regressor proposed in [5]. Within each subplot, from left to right, the performance of continual learning models using HM (green), MM (yellow), and LM (blue) images for replays is presented in Acc (left), BWT (middle), and Forgetting (right). See Sec. 4.3 for the introduction to the evaluation metrics and the continual learning models. The error bars are the standard errors calculated over 5 independent runs with different random seeds.
A.5 Performance of Models Using ICMscore without Time Penalty
We modify ICMscore by removing the time penalty, defining $s_{i}^{\prime}=1$ (correct) and $s_{i}^{\prime}=0$ (incorrect), while keeping ICMscore(I)$=\frac{\sum_{i=1}^{n}s_{i}}{n}$. Following the dataset splitting protocol from Sec. 4.2, we train image recognition and continual learning models on LM, MM, and HM subsets.
Figure A6: Performance of Models Using ICMscore without time penalty. (a) Top-1 accuracy comparison for image classification models trained on LM (blue), HM (green), and MM (yellow) images. The dashed line indicates the 20% chance level. (b) Continual learning performance showing Accuracy (left), BWT (middle), and Forgetting (right) for models using HM (green), MM (yellow), and LM (blue) replays. Error bars represent standard error over 5 independent runs. Evaluation metrics are detailed in Sec. 4.3.

### v5

@@ E.2 Performance of Different Continual Learning Models
Figure A6: Continual learning performance of AI models using HM, MM, and LM images for replays. (a) shows results using our ICMscore predictor from Sec. 4.1, while (b) presents results from the regressor proposed in Khosla et al. (2015). Within each subplot, from left to right, the performance of continual learning models using HM (green), $M$ (yellow), and LM (blue) images for replays is presented in Acc (left), BWT (middle), and Forgetting (right). See Sec. 4.3 for the introduction to the evaluation metrics and the continual learning models. The error bars are the standard errors calculated over 5 independent runs with different random seeds.
Similar to the image recognition experiments in Sec. 4.2, we categorize images into HM, MM, and LM using both our ICMscore predictor and the predictor from Khosla et al. (2015). From Fig. A6, replay methods based on the predictor from Khosla et al. (2015) show no significant performance difference between LM and HM images. In contrast, our ICMscore predictor clearly reveals substantial differences: continual learning models replaying LM images consistently outperform those replaying HM images, as well as models using replay images selected by the predictor from Khosla et al. (2015). See Sec. 4.3 for a detailed discussion.
Beyond such naive replay strategies, we also evaluate more recent continual learning approaches. In particular, we adopt Learning to Prompt (L2P) Wang et al. (2022), which employs a small pool of key-value embeddings, or prompts, to encode task-specific knowledge while leveraging a frozen pretrained backbone. The input image feature representation is matched against the prompt keys via cosine similarity, and the top-$k$ prompts are prepended to the input embedding for downstream classification. To adapt this framework to our study, we extend L2P with a replay buffer that stores a random subset of previously seen examples. This design emphasizes the influence of memorability-aware replay data while disentangling its role during training. For the backbone, instead of the official ImageNet-pretrained ViT, we use a DINO-pretrained ViT Caron et al. (2021) to mitigate label leakage from ImageNet, ensuring a cleaner evaluation.
Following the same continual learning setup as in Sec. 4.3, we structure ICMD into 6 tasks by first training on 5 object classes and then sequentially introducing 1 incremental class at a time. Three variants of the augmented L2P model are trained, each using only HM, MM, or LM images for replay, while keeping the training images constant across variants. All models are trained for 10 epochs in the initial task and 5 epochs for incremental tasks, using a constant learning rate of 0.003. Other hyperparameter settings follow the original L2P configuration.
((a))
((b))
Figure A7: Continual learning performance of Learning to Prompt (L2P) Wang et al. (2022) augmented with memorability-based replay buffers. (a) Performance comparison of L2P models that stores HM (green), $M$ (yellow), and LM (blue) images for the additional replay buffer, evaluated across Accuracy (left), BWT (middle), and Forgetting (right) metrics. Error bars represent standard errors calculated over 5 independent runs with different random seeds. (b) Task progression showing Accuracy at last task (solid lines) and Forgetting (dotted lines) as functions of task number for L2P models with HM (green), $M$ (yellow), and LM (blue) images stored in the replay buffer.
Results in Fig. A7 show that L2P models using LM images for replay consistently outperform those using HM and MM images across all three metrics (Acc, BWT, and F), which aligns with the patterns observed in naive replay methods.
Finally, following the memorability predictor protocol in Sec. 4.3, we repeat the L2P experiments using both our ICMscore predictor and the regressor from Khosla et al. (2015). As shown in Fig. A8, the latter yields minimal differences between LM and HM replay, whereas our ICMscore predictor produces clear and consistent performance gaps. These results further highlight the advantages of memorability-based selection strategies.
((a))
((b))
Figure A8: Continual learning performance of Learning to Prompt (L2P) Wang et al. (2022) augmented with memorability-based replay buffers using different predictors. (a) Results with our ICMscore predictor from Sec. 4.1. (b) Results with the regressor from Khosla et al. (2015). Metrics reported are Accuracy (left), BWT (middle), and Forgetting (right). Error bars indicate standard error over 5 independent runs with different random seeds.